# publications.bib

### v3

@@ 4.1. The Spectral Action
    $\mathrm{WD}_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}H(\overline{\mathbb{Q}}_{\ell})$
    where $\overline{\mathbb{Q}}_{\ell}$ has the discrete topology, $\mathrm{SL}(2,\overline{\mathbb{Q}}_{\ell})$ acts via an algebraic representation, and the map respects the action of $W_{\mathbb{Q}_{p}}$ on $\phantom{}{}^{L}H(\overline{\mathbb{Q}}_{\ell})$, the $L$-group of $H$,
  • the set $\Phi^{\mathrm{ss}}(H)$ of continuous semi-simple homomorphisms
@@
Then, we assume, for all $b\in B(G)$, that there exists a map
@@
$\rho\mapsto\phi_{\rho}$
satisfying the following properties:
@@
[…]
@@
    Consider $\phi_{\rho}$ as an element of $\Phi(G)$ given by composing with the twisted embedding $\phantom{}{}^{L}J_{b}(\overline{\mathbb{Q}}_{\ell})\simeq\phantom{}^{L}M_{b}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}G(\overline{\mathbb{Q}}_{\ell})$ (as defined in [FS, Section IX.7.1]). Then $\phi_{\rho}$ factors through the natural embedding $\phantom{}{}^{L}T\rightarrow\phantom{}^{L}G$ if and only if $b\in B(G)_{\mathrm{un}}$.
@@
    If $\rho$ is a representation such that $W_{\mathbb{Q}_{p}}\times\mathrm{SL}(2,\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}J_{b}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}G(\overline{\mathbb{Q}}_{\ell})$ factors through $\phantom{}{}^{L}T$, where the last map is the twisted embedding then, by (2), the element $b$ is unramified, and we require that $\rho$ is isomorphic to an irreducible constituent of $\rho_{b,w}$ for $w\in W_{b}$ and $\chi$ the character attached to the induced toral parameter $\phi_{T}$.
@@
Assuming 4.4, we have that the following is true for a parameter $\phi$ induced from a generic parameter $\phi_{T}$. Given any Schur-irreducible object $A\in\mathrm{D}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{F}}_{\ell})_{\phi}\subset\mathrm{D}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{F}}_{\ell})\simeq\mathrm{D}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$ then $A$ is non-zero if and only if $b\in B(G)_{\mathrm{un}}$, and in this case it must be an irreducible sub-quotient of $\rho_{b,w}$, for some $w\in W_{b}$.
@@
This follows from combining the proof of [Ham2, Corollary 7.7] with Lemma 4.2 (1). ∎
Since we want some flexibility in the groups for which we have the above results, we discuss how Assumption 4.4 behaves under central isogenies.
@@
We consider an injective map $\psi:G^{\prime}\hookrightarrow G$ of connected reductive groups which induces an isomorphism of adjoint and derived groups, and the induced map $\psi_{\mathrm{Bun}}:B(G^{\prime})\rightarrow B(G)$ on the associated Kottwitz sets. We now have the following lemma.
@@
If $\psi:G^{\prime}\rightarrow G$ is an injective map which induces an isomorphism on adjoint and derived groups then it follows that $\psi_{\mathrm{Bun}}:B(G^{\prime})\rightarrow B(G)$ induces an injection $J_{b^{\prime}}\rightarrow J_{b}$ which is an isomorphism of the derived group and adjoint groups for all $b=\psi_{\mathrm{Bun}}(b^{\prime})$ and $b^{\prime}\in B(G^{\prime})$.
@@
Since $\psi$ is an inclusion it easily follows that it induces an inclusion $J_{b^{\prime}}\rightarrow J_{b}$ of $\sigma$-centralizers. To see that it induces an isomorphism on derived/adjoint groups, recall that $J_{b}$ is an inner form of a Levi subgroup $M_{b}$ of $G$ given by the centralizer of the slope homomorphism of $b$. The preimage of $M_{b}$ under $\phi$ defines a Levi subgroup $M_{b^{\prime}}$ of $G$ which will be the centralizer of the slope homomorphism of $b^{\prime}$, since $\phi$ induces an isomorphism on adjoint groups. Moreover, the inner twisting from $J_{b}$ to $M_{b}$ and $J_{b^{\prime}}$ to $M_{b^{\prime}}$ are compatible with the inclusion in the sense that the inclusion $J_{b^{\prime}}\rightarrow J_{b}$ is given by applying the inner twist of $M_{b}$ to the inclusion $M_{b^{\prime}}\rightarrow M_{b}$. Since the formation of derived/adjoint groups respects inner twists, this reduces us to showing that the map $M_{b^{\prime}}\rightarrow M_{b}$ on Levi subgroups induces an isomorphism on the derived/adjoint groups, and this is clear. ∎
@@
[…]
commute, where $b:=\psi_{\mathrm{Bun}}(b^{\prime})$. Here the right vertical arrow is given by composing a parameter $\phi:\mathrm{WD}_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}J_{b}(\overline{\mathbb{Q}}_{\ell})$ with the induced map $\phantom{}{}^{L}J_{b}\rightarrow\phantom{}^{L}J_{b^{\prime}}$ on the dual groups, and the left vertical arrow is not a map at all, it is a correspondence defined by the subset of $\Pi(J_{b})\times\Pi(J_{b^{\prime}})$ consisting of pairs $(\pi_{b},\pi_{b^{\prime}})$ such that $\pi_{b^{\prime}}$ is a constituent of the restriction of $\pi_{b}$ to $J_{b^{\prime}}(\mathbb{Q}_{p})$. We will now show that this gives rise to a well-defined map under our assumptions on $\psi$. Given a representation $\pi_{b^{\prime}}\in\Pi(J_{b^{\prime}})$, it follows by [GK, Lemma 2.3] and the previous Lemma that we can find a lift $\pi_{b}\in\Pi(J_{b})$ such that $\pi_{b^{\prime}}$ is an irreducible constituent of $\pi_{b}|_{J_{b^{\prime}}(\mathbb{Q}_{p})}$. It also follows from [GK, Lemma 2.1] and [Tad1, Proposition 2.4,Corollary 2.5] that the set $\Pi_{\pi_{b}}(J_{b^{\prime}})$ of representations of $J_{b^{\prime}}$ occurring in the restriction of $\pi_{b}$ is finite. Now, using the previous Lemma, we have the following.
@@
[GK, Lemma 2.4] For the map $J_{b^{\prime}}\rightarrow J_{b}$ of $\sigma$-centralizers induced by a map $\psi$ as above, and $\pi_{b}^{1},\pi_{b}^{2}\in\Pi(J_{b})$ the following are equivalent.
@@
Now we can use this to define $\mathrm{LLC}_{b^{\prime}}:\Pi(J_{b^{\prime}})\rightarrow\Phi(J_{b^{\prime}})$ in terms of $\mathrm{LLC}_{b}:\Pi(J_{b})\rightarrow\Phi(J_{b})$ for $\mathrm{LLC}_{\mathrm{Bun}_{G}}$ satisfying Assumption 4.4. Namely, for $\pi_{b^{\prime}}\in\Pi(J_{b^{\prime}})$, we let $\pi_{b}\in\Pi(J_{b})$ be a representation such that $\pi_{b^{\prime}}$ occurs as an irreducible constituent of $\pi_{b}|_{J_{b^{\prime}}(\mathbb{Q}_{p})}$. We set $\phi_{\pi_{b^{\prime}}}$ to be the parameter $\phi_{\pi_{b}}$ attached to $\pi_{b}$ under $\mathrm{LLC}_{b}$ composed with the map $\phantom{}{}^{L}J_{b}\rightarrow\phantom{}^{L}J_{b^{\prime}}$ on dual groups induced by $\psi$. By the previous Lemma, any two choices of lifts $\pi^{1}_{b}$ and $\pi^{2}_{b}$ of $\pi_{b^{\prime}}$ will differ by a character twist of $\chi\in(J_{b}(\mathbb{Q}_{p})/J_{b^{\prime}}(\mathbb{Q}_{p}))^{\vee}$. We note that the Fargues-Scholze local Langlands correspondence is compatible with character twists [FS, Theorem I.9.6 (ii)]. Since $\mathrm{LLC}_{b}$ is compatible with the Fargues-Scholze local Langlands after semi-simplification by assumption, it follows that the same is true for $\mathrm{LLC}_{b}$. Therefore, $\phi_{\pi^{1}_{b}}$ and $\phi_{\pi^{2}_{b}}$ differ by a character twist that becomes trivial after composing with $\phantom{}{}^{L}J_{b}\rightarrow\phantom{}^{L}J_{b^{\prime}}$, and so $\phi_{\pi_{b}}$ does not depend on the choice of lift. We let $\mathrm{LLC}_{\mathrm{Bun}_{G^{\prime}}}:\bigsqcup_{b^{\prime}\in B(G^{\prime})}\Pi(J_{b^{\prime}})\rightarrow\Phi(G)$ be the local Langlands correspondence defined by the $\mathrm{LLC}_{b^{\prime}}$ for $b^{\prime}$ varying. We now prove that our assumption is comaptible with central isogenies.
@@
Suppose we have an injective map $\psi:G^{\prime}\hookrightarrow G$ of quasi-split connected reductive groups inducing an isomorphism on adjoint and derived groups. Assume we have a local Langlands correspondence $\mathrm{LLC}_{\mathrm{Bun}_{G}}:\bigsqcup_{b\in B(G)}\Pi(J_{b})\rightarrow\Phi(G)$ such that Assumption 4.4 holds. If we let $\mathrm{LLC}_{\mathrm{Bun}_{G^{\prime}}}:\bigsqcup_{b^{\prime}\in B(G^{\prime})}\Pi(J_{b^{\prime}})\rightarrow\Phi(G^{\prime})$ be the local Langlands correspondence induced by $\mathrm{LLC}_{\mathrm{Bun}_{G}}$ and $\psi$ as above then $\mathrm{LLC}_{\mathrm{Bun}_{G^{\prime}}}$ satisfies Assumption 4.4 as well.
@@
We note, since the Fargues-Scholze local Langlands correspondence is compatible with maps $G^{\prime}\rightarrow G$ that induce an isomorphism of adjoint groups [FS, Theorem I.9.6 (v)], it follows by the above construction that if Assumption 4.4 (1) holds true for $\mathrm{LLC}_{\mathrm{Bun}_{G}}$ then it also holds true for $\mathrm{LLC}_{\mathrm{Bun}_{G^{\prime}}}$. Suppose we have $b^{\prime}\in B(G^{\prime})$ mapping to $b\in B(G)_{\mathrm{un}}$. We let $B_{b}\subset J_{b}$ be the corresponding Borel then, since the map $J_{b^{\prime}}\rightarrow J_{b}$ induces an isomorphism on adjoint groups by Lemma 4.6, it follows that $B\cap J_{b^{\prime}}=:B_{b^{\prime}}\subset J_{b^{\prime}}$ is a Borel of $J_{b^{\prime}}$. In particular, $b^{\prime}$ must be an unramified element of $B(G^{\prime})_{\mathrm{un}}$. Now, the map $J_{b^{\prime}}\rightarrow J_{b}$ induces an isomorphism
@@
If we let $T$ be the maximal split torus of $J_{b}$ then the preimage $T^{\prime}$ under $\phi$ is a maximal torus of $J_{b^{\prime}}$, and the previous isomorphism of flag varieties implies that, given a character $\chi:T(\mathbb{Q}_{p})\rightarrow\overline{\mathbb{Q}}_{\ell}^{*}$, we have an isomorphism:
@@
$\mathrm{WD}_{\mathbb{Q}_{p}}\xrightarrow{\phi_{\pi_{b}}}\phantom{}^{L}J_{b}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}J_{b^{\prime}}(\overline{\mathbb{Q}}_{\ell})$
@@
Now that we have shown this compatibility assumption is somewhat flexible, we can state the groups we know to satisfy Assumption 4.4. This result is largely contained in [Ham1, FS, HKW, BMNH], but we also want to consider an additional group $\mathrm{GU}_{2}$, where we have the following construction of $\mathrm{LLC}_{\mathrm{Bun}_{\mathrm{GU}_{2}}}$. Recall that $\mathrm{GU}_{2}/L$ can be written as
@@
There is a bijection between $\Pi(J_{b^{\prime}})$ and the set of pairs $(\tilde{\pi},\chi)$ such that $\tilde{\pi}\in\Pi(J_{\tilde{b}})$ and $\chi$ is a character of $(L^{\prime})^{\times}$ such that $\chi|_{L^{\times}}=\omega_{\tilde{\pi}}|_{L^{\times}}$, where $\omega_{\tilde{\pi}}$ is the central character of $\tilde{\pi}$.
@@
of groups over $L$. In particular, we see that $J_{b}=J_{\tilde{b}}\times\mathrm{Res}_{L^{\prime}/L}\mathbb{G}_{m}$, and the quotient map $J_{b}\rightarrow J_{b^{\prime}}$ induces an isomorphism on adjoint and derived subgroups of $J_{b^{\prime}}$. Moreover, by Hilbert’s Theorem 90, we have $H^{1}(L,\mathbb{G}_{m})=0$, and thus we also have a surjection on $L$-points, from which the lemma follows. To see this isomorphism, recall that $J_{b}$ (resp. $J_{b^{\prime}},J_{\tilde{b}}$) is an inner form of $M_{b}$ (resp. $M_{b^{\prime}},M_{\tilde{b}}$), the Levi subgroup of $H$ (resp. $\mathrm{GU}_{2},\mathrm{GL}_{2}$) given by the centralizer of the slope homomorphism of $b$ (resp. $b^{\prime},\tilde{b}$). In particular, we see that we have an isomorphism

### v5

@@ 4.2.1. Implications of Compatibility for Generic Parameters
    $W_{\mathbb{Q}_{p}}\times\mathrm{SL}(2,\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}H(\overline{\mathbb{Q}}_{\ell})$
    where $\overline{\mathbb{Q}}_{\ell}$ has the discrete topology, $\mathrm{SL}(2,\overline{\mathbb{Q}}_{\ell})$ acts via an algebraic representation, $W_{\mathbb{Q}_{p}}$ acts through a semi-simple homorphism (in the sense that if it factors through $\phantom{}{}^{L}P$ for a proper parabolic $P\subset G$ then it also factors through the associated Levi $\phantom{}{}^{L}M$), and the map respects the action of $W_{\mathbb{Q}_{p}}$ on $\phantom{}{}^{L}H(\overline{\mathbb{Q}}_{\ell})$, the $L$-group of $H$,
  • the set $\Phi^{\mathrm{ss}}(H)$ of continuous (where again $\overline{\mathbb{Q}}_{\ell}$ has the discrete topology) semi-simple homomorphisms,
@@
Then, we assume that there exists a collection of maps
@@
$\rho\mapsto\phi_{\rho},$
for all $b\in B(G)$, satisfying the following properties:
@@
    $\Pi(J_{b})$$\Phi(J_{b})$$\Phi^{\mathrm{ss}}(J_{b})$$\scriptstyle{\mathrm{LLC}_{b}}$$\scriptstyle{\mathrm{LLC}^{\mathrm{FS}}_{b}}$$\scriptstyle{(-)^{\mathrm{ss}}}$
@@
    For $\rho$ a non-zero smooth irreducible representation of $J_{b}(\mathbb{Q}_{p})$ for some $b\in B(G)$, consider $\phi_{\rho}$ as an element of $\Phi(G)$ given by composing with the twisted embedding $\phantom{}{}^{L}J_{b}(\overline{\mathbb{Q}}_{\ell})\simeq\phantom{}^{L}M_{b}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}G(\overline{\mathbb{Q}}_{\ell})$ (as defined in [16, Section IX.7.1]). Then $\phi_{\rho}$ factors through the natural (up to $\hat{G}-$conjugacy) embedding $\phantom{}{}^{L}T\rightarrow\phantom{}^{L}G$ if and only if $b\in B(G)_{\mathrm{un}}$.
@@
    If $\rho$ is a non-zero representation of $J_{b}(\mathbb{Q}_{p})$ such that $W_{\mathbb{Q}_{p}}\times\mathrm{SL}(2,\overline{\mathbb{Q}}_{\ell})\xrightarrow{\phi_{\rho}}\phantom{}^{L}J_{b}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}G(\overline{\mathbb{Q}}_{\ell})$ factors through the embedding $\phantom{}{}^{L}T\rightarrow\phantom{}^{L}G$ with induced parameter $\phi_{T}$, then, by (2), the element $b$ is unramified, and we require that $\rho$ is isomorphic to an irreducible subquotient of $i_{B_{b}}^{J_{b}}(\chi^{w})\otimes\delta_{P_{b}}^{-1/2}$ for $w\in W_{b}$, where $\chi:T(\mathbb{Q}_{p})\rightarrow\overline{\mathbb{Q}}_{\ell}^{*}$ is the character attached to $\phi_{T}$ by class field theory and the notation is as above. Here the map $\phantom{}{}^{L}T\rightarrow\phantom{}^{L}G$ is the non-twisted embedding and the map $\phantom{}{}^{L}J_{b}\rightarrow\phantom{}^{L}G$ is the twisted embedding.
@@
Assuming 4.5, we have that the following is true for a parameter $\phi$ induced from a generic parameter $\phi_{T}$. Given any Schur-irreducible object $A\in\mathrm{D}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{F}}_{\ell})_{\phi}\subset\mathrm{D}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{F}}_{\ell})\simeq\mathrm{D}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$ then $A$ is non-zero if and only if $b\in B(G)_{\mathrm{un}}$, and in this case it must be an irreducible sub-quotient of $\rho_{b,w}$, for some $w\in W_{b}$.
@@
This follows from combining the proof of [24, Corollary 6.7] with Lemma 4.3 (1). ∎
Since we want some flexibility in the groups for which we have the above results, we discuss how Assumption 4.5 behaves under central isogenies, and then spell out the main relevant cases for which this assumption holds.
@@
We consider an injective map $\psi:G^{\prime}\hookrightarrow G$ of connected reductive groups which induces an isomorphism of derived groups, and the induced map $\psi_{\mathrm{Bun}}:B(G^{\prime})\rightarrow B(G)$ on the associated Kottwitz sets. We now have the following Lemma.
@@
If $\psi:G^{\prime}\rightarrow G$ is an injective map which induces an isomorphism on derived groups then it follows that $\psi_{\mathrm{Bun}}:B(G^{\prime})\rightarrow B(G)$ induces an injection $\psi_{b^{\prime}}:J_{b^{\prime}}\rightarrow J_{b}$ for all $b=\psi_{\mathrm{Bun}}(b^{\prime})$ and $b^{\prime}\in B(G^{\prime})$, which is an isomorphism on the derived groups.
@@
Since $\psi$ is an inclusion it easily follows that it induces an inclusion $J_{b^{\prime}}\rightarrow J_{b}$ of $\sigma$-centralizers. To see that it induces an isomorphism on derived groups, recall that $J_{b}$ is an inner form of a Levi subgroup $M_{b}$ of $G$ given by the centralizer of the slope homomorphism of $b$. The preimage of $M_{b}$ under $\psi$ defines a Levi subgroup $M_{b^{\prime}}$ of $G$ which will be the centralizer of the slope homomorphism of $b^{\prime}$, since $\psi$ induces an isomorphism on derived groups and in turn also on adjoint groups (since $(G^{\mathrm{der}})^{\mathrm{ad}}\simeq G^{\mathrm{ad}}$)). Moreover, the inner twisting from $J_{b}$ to $M_{b}$ and $J_{b^{\prime}}$ to $M_{b^{\prime}}$ are compatible with the inclusion in the sense that the inclusion $J_{b^{\prime}}\rightarrow J_{b}$ is given by applying the inner twist of $M_{b}$ to the inclusion $M_{b^{\prime}}\rightarrow M_{b}$. Since the formation of derived groups respects inner twists, this reduces us to showing that the map $M_{b^{\prime}}\rightarrow M_{b}$ on Levi subgroups induces an isomorphism on the derived groups, and this is clear. ∎
@@
$\Pi(J_{b})$$\Phi(J_{b})$$\Pi(J_{b^{\prime}})$$\Phi(J_{b^{\prime}})$$\scriptstyle{\mathrm{LLC}_{b}}$$\scriptstyle{\mathrm{LLC}_{b^{\prime}}}$
commute, where $b:=\psi_{\mathrm{Bun}}(b^{\prime})$. Here the right vertical arrow is given by composing a parameter $\phi:W_{\mathbb{Q}_{p}}\times\mathrm{SL}_{2}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}J_{b}(\overline{\mathbb{Q}}_{\ell})$ with the induced map $\phantom{}{}^{L}J_{b}\rightarrow\phantom{}^{L}J_{b^{\prime}}$ on the dual groups, and the left vertical arrow is not a map at all, it is a correspondence defined by the subset of $\Pi(J_{b})\times\Pi(J_{b^{\prime}})$ consisting of pairs $(\pi_{b},\pi_{b^{\prime}})$ such that $\pi_{b^{\prime}}$ is a constituent of the restriction of $\pi_{b}$ to $J_{b^{\prime}}(\mathbb{Q}_{p})$. We will now show that this gives rise to a well-defined map under our assumptions on $\psi$. Given a representation $\pi_{b^{\prime}}\in\Pi(J_{b^{\prime}})$, it follows by [63, Proposition 2.2] and the previous Lemma that we can find a lift $\pi_{b}\in\Pi(J_{b})$ such that $\pi_{b^{\prime}}$ is an irreducible constituent of $\pi_{b}|_{J_{b^{\prime}}(\mathbb{Q}_{p})}$. It also follows from [62, Lemma 2.1] that the set $\Pi_{\pi_{b}}(J_{b^{\prime}})$ of representations of $J_{b^{\prime}}$ occurring in the restriction of $\pi_{b}$ is finite. Now, using the previous Lemma, we have the following.
@@
[21, Corollary 2.5] For the map $J_{b^{\prime}}\rightarrow J_{b}$ of $\sigma$-centralizers induced by a map $\psi$ as above and $\pi_{b}^{1},\pi_{b}^{2}\in\Pi(J_{b})$, the following are equivalent.
@@
Now we can use this to define $\mathrm{LLC}_{b^{\prime}}:\Pi(J_{b^{\prime}})\rightarrow\Phi(J_{b^{\prime}})$ in terms of $\mathrm{LLC}_{b}:\Pi(J_{b})\rightarrow\Phi(J_{b})$ for $\mathrm{LLC}_{\mathrm{Bun}_{G}}$ satisfying Assumption 4.5. Namely, for $\pi_{b^{\prime}}\in\Pi(J_{b^{\prime}})$, we let $\pi_{b}\in\Pi(J_{b})$ be a representation such that $\pi_{b^{\prime}}$ occurs as an irreducible constituent of $\pi_{b}|_{J_{b^{\prime}}(\mathbb{Q}_{p})}$. We set $\phi_{\pi_{b^{\prime}}}$ to be the parameter $\phi_{\pi_{b}}$ attached to $\pi_{b}$ under $\mathrm{LLC}_{b}$ composed with the map $\phantom{}{}^{L}J_{b}\rightarrow\phantom{}^{L}J_{b^{\prime}}$ on dual groups induced by $\psi$. By the previous Lemma, any two choices of lifts $\pi^{1}_{b}$ and $\pi^{2}_{b}$ of $\pi_{b^{\prime}}$ will differ by a character twist of $\chi\in(J_{b}(\mathbb{Q}_{p})/J_{b^{\prime}}(\mathbb{Q}_{p}))^{\vee}$. We note that the Fargues-Scholze local Langlands correspondence is compatible with character twists [16, Theorem I.9.6 (ii)]. Since $\mathrm{LLC}_{b}$ is compatible with the Fargues-Scholze local Langlands after semi-simplification by assumption, it follows that the same is true for $\mathrm{LLC}_{b^{\prime}}$. Therefore, $\phi_{\pi^{1}_{b}}$ and $\phi_{\pi^{2}_{b}}$ differ by a character twist that becomes trivial after composing with $\phantom{}{}^{L}J_{b}\rightarrow\phantom{}^{L}J_{b^{\prime}}$, and so $\phi_{\pi_{b}}$ does not depend on the choice of lift. We let $\mathrm{LLC}_{\mathrm{Bun}_{G^{\prime}}}:\bigsqcup_{b^{\prime}\in B(G^{\prime})}\Pi(J_{b^{\prime}})\rightarrow\Phi(G)$ be the local Langlands correspondence defined by the $\mathrm{LLC}_{b^{\prime}}$ for $b^{\prime}$ varying. We now prove that our assumption is compatible with central isogenies.
@@
Suppose we have an injective map $\psi:G^{\prime}\hookrightarrow G$ of quasi-split connected reductive groups inducing an isomorphism on derived groups. Assume we have a local Langlands correspondence $\mathrm{LLC}_{\mathrm{Bun}_{G}}:\bigsqcup_{b\in B(G)}\Pi(J_{b})\rightarrow\Phi(G)$ such that Assumption 4.5 holds. If we let $\mathrm{LLC}_{\mathrm{Bun}_{G^{\prime}}}:\bigsqcup_{b^{\prime}\in B(G^{\prime})}\Pi(J_{b^{\prime}})\rightarrow\Phi(G^{\prime})$ be the local Langlands correspondence induced by $\mathrm{LLC}_{\mathrm{Bun}_{G}}$ and $\psi$ as above then $\mathrm{LLC}_{\mathrm{Bun}_{G^{\prime}}}$ satisfies Assumption 4.5 as well.
@@
We note, since the Fargues-Scholze local Langlands correspondence is compatible with maps $G^{\prime}\rightarrow G$ that induce an isomorphism of adjoint groups [16, Theorem I.9.6 (v)] (which follows from the map inducing an isomorphism on derived groups), it follows by the above construction that if Assumption 4.5 (1) holds true for $\mathrm{LLC}_{\mathrm{Bun}_{G}}$ then it also holds true for $\mathrm{LLC}_{\mathrm{Bun}_{G^{\prime}}}$. Suppose we have $b^{\prime}\in B(G^{\prime})$ mapping to $b\in B(G)_{\mathrm{un}}$. We let $B_{b}\subset J_{b}$ be the corresponding Borel then, since the map $J_{b^{\prime}}\rightarrow J_{b}$ induces an isomorphism on adjoint groups by Lemma 4.7, it follows that $B\cap J_{b^{\prime}}=:B_{b^{\prime}}\subset J_{b^{\prime}}$ is a Borel of $J_{b^{\prime}}$. In particular, $b^{\prime}$ must be an unramified element of $B(G^{\prime})_{\mathrm{un}}$. Now, the map $J_{b^{\prime}}\rightarrow J_{b}$ induces an isomorphism
@@
If we let $T$ be the maximal split torus of $J_{b}$ then the preimage $T^{\prime}$ under $\psi_{b^{\prime}}$ the map induced by $\psi$ is a maximal torus of $J_{b^{\prime}}$, and the previous isomorphism of flag varieties implies that, given a character $\chi:T(\mathbb{Q}_{p})\rightarrow\overline{\mathbb{Q}}_{\ell}^{*}$, we have an isomorphism:
@@
Indeed, recall that the left hand side is described by the set of compactly supported smooth functions on $J_{b}(\mathbb{Q}_{p})$ which transform under $B_{b}(\mathbb{Q}_{p})$ via the character $\chi$ of $T(\mathbb{Q}_{p})$ inflated to $B_{b}(\mathbb{Q}_{p})$. In particular, by the previous isomorphism of flag varieties, restricting to $J_{b^{\prime}}(\mathbb{Q}_{p})$ gives the collection of compactly supported smooth (since $J_{b^{\prime}}(\mathbb{Q}_{p})\rightarrow J_{b}(\mathbb{Q}_{p})$ will be a closed subgroup, by Lemma 4.7) functions on $J_{b^{\prime}}(\mathbb{Q}_{p})$ which transform under $B_{b^{\prime}}(\mathbb{Q}_{p})$ by the restriction of the inflated character $\chi$ to $B_{b^{\prime}}(\mathbb{Q}_{p})$, but this is just the inflation of $\chi|_{T^{\prime}(\mathbb{Q}_{p})}$ to $B_{b^{\prime}}(\mathbb{Q}_{p})$, giving the desired claim.
@@
$W_{\mathbb{Q}_{p}}\times\mathrm{SL}_{2}(\overline{\mathbb{Q}}_{\ell})\xrightarrow{\phi_{\pi_{b}}}\phantom{}^{L}J_{b}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}J_{b^{\prime}}(\overline{\mathbb{Q}}_{\ell})$
@@
Now that we have shown this compatibility assumption is somewhat flexible, we can state the groups we know to satisfy Assumption 4.5. This result is largely contained in [27, 16, 31, 47], but we also want to consider an additional group $\mathrm{GU}_{2}$, where we have the following construction of $\mathrm{LLC}_{\mathrm{Bun}_{\mathrm{GU}_{2}}}$. For $L/\mathbb{Q}_{p}$ a finite extension, recall that $\mathrm{GU}_{2}/L$ can be written as
@@
[…]
@@
of groups over $L$. Assuming this, we see that since by Hilbert’s Theorem 90 we have $H^{1}(L,\mathbb{G}_{m})=0$, the projection map from $(J_{\tilde{b}}\times\mathrm{Res}_{L^{\prime}/L}\mathbb{G}_{m})$ to $J_{b^{\prime}}$ induces a short exact sequence on $L$-points
$0\rightarrow L^{\times}\rightarrow J_{\tilde{b}}(L)\times(L^{\prime})^{\times}\rightarrow J_{b^{\prime}}(L)\rightarrow 0,$
so we see that to give a smooth irreducible $\overline{\mathbb{Q}}_{\ell}$-representation of $J_{b^{\prime}}(L)$ is exactly equivalent to giving a smooth irreducible $\overline{\mathbb{Q}}_{\ell}$-representation of $J_{\tilde{b}}(L)$ and $(L^{\prime})^{\times}$ which agree on the restriction to the central copy of $L^{\times}$.
To see this isomorphism, recall that $J_{b}$ (resp. $J_{b^{\prime}},J_{\tilde{b}}$) is an inner form of $M_{b}$ (resp. $M_{b^{\prime}},M_{\tilde{b}}$), the Levi subgroup of $H$ (resp. $\mathrm{GU}_{2},\mathrm{GL}_{2}$) given by the centralizer of the slope homomorphism of $b$ (resp. $b^{\prime},\tilde{b}$). Note that we have $J_{b}=J_{\tilde{b}}\times\mathrm{Res}_{L^{\prime}/L}\mathbb{G}_{m}\subset H$, and the quotient map $J_{b}\rightarrow J_{b^{\prime}}$ induces an isomorphism on adjoint and derived subgroups of $J_{b^{\prime}}$. Moreover, we see that we have an isomorphism